A piston of radins R=20 cm is free to move without friction inside a vertical cylinder filled with water and having the same diameter as that of the piston. The piston that floats at the surface of water is traversed by a hollow vertical straw with a radius r =1 cm Caleulate the height of water in the straw if the piston weighs 1250 N. P-1000 Ka/m3


1
Expert's answer
2021-06-07T19:16:17-0400

"p=\frac{F}{S_1}=\frac{\rho Vg}{S_2}=\frac{\rho h S_1 g}{S_2},\implies"

"h=\frac{FS_2}{\rho S_1^2 g}=\frac{F}{\rho g}\cdot (\frac{r}{R^2})^2=12.5~cm."